The planet Mercury's surface temperature varies from 700 K during the day to 90 K at night.What is the temperature in ?C during the day?What is the temperature in ?C at night?What is the temperature in ?F during the day?What is the temperature in ?F at night?
Answer #1

a)

Kelvin to Degree centigrade

Tday=700-273 = 427oC

b)

Tnight=90-273 =-183oC

c)

Degrees to fahrenheit

oF=oC*(9/5) +32

so

Tday=427*(9/5) +32=800.6 oF

d)

Tnight=-183*(9/5)+32 =-297.4 oF